Fifteen years ago, Hailey invested $5,000 and locked in an annual interest rate of 6 percent for 30 years (ending 15 years from now). Aidan can make a 15-year investment today and lock in an interest rate of 8 percent.

How much money should he invest now in order to have the same amount of money in 15 years as Hailey?
Answer #1

AMount Required after 15 Years from Now = Hailey's amount

= CF * ( 1+r)^n

= $ 5,000 * ( 1+ 0.06)^ 30

= $ 5000 (1.06)^30

= $ 5000 * 5.7435

= 28717.46

AMount to be deposited by Aiden =

= CF ( 1+0.08) ^ 15

= CF ( 1.08)^15

= CF * 3.1722

28717.46 = CF * 3.1722

CF = 28717.46 / 3.1722

= $ 9052.94